# Bridging doubly heavy tetraquark mass spectrum with heavy baryons utilizing heavy antiquark-diquark symmetry

### v1

@@ II.2 Mass formula of DHT in CQM
After determining the mass of the heavy diquark $m_{[QQ^{\prime}]^{C}_{S}}$, we need the binding and spin hyperfine coupling constant between $QQ^{\prime}]^{C}_{S}$} and the light quark to calculate the masses of DHTs with Eq. (5). In Refs. Karliner et al. (2009); Karliner and Rosner (2014); Wu and Ma (2023), the studies indicate that there are no bindings between light quark pair ($qq^{\prime}$) and heavy-unflavored quark pair $Qn$. Since the color and spin operators have been explicitly introduced in Eq. (5), the binding between heavy diquark and light quark should be independent of both color and spin, denoted as $B_{[QQ^{\prime}]q}$. So we have $B_{[QQ^{\prime}]q}=B_{Qq}+B_{Q^{\prime}q}$, resulting in $B_{[QQ^{\prime}]n}=0$, $B_{[cc]s}=2B_{cs}=106.8$ MeV, $B_{[bb]s}=2B_{bs}=125.2$ MeV, and $B_{[cb]s}=B_{cs}+B_{bs}=116.0$ MeV. For the spin hyperfine coupling constant $\alpha_{ij}$, the constant is the same for the light-light quarks and heavy-light quarks but is flavor dependent for the heavy-heavy quarks. For $\alpha_{[QQ^{\prime}]^{C}_{S}q}$, since the heavy diquark is regarded as a heavy quark $X$ with HADS, $\alpha_{[QQ^{\prime}]^{C}_{S}q}$ should be $\alpha_{Xq}$, which is proved to be independent on the mass of heavy quark $X$ and share the same value as $\alpha_{qq^{\prime}}$ (abbreviated as $\alpha$) . For detailed discussions, refer to Refs. Karliner et al. (2009); Karliner and Rosner (2014); Wu and Ma (2023). The masses, bindings and spin hyperfine splittings ($\alpha_{ij}/m_{i}m_{j}$) needed for calculating the masses of DHTs are listed in Table 2. It should be noted that no new parameters are introduced here. All parameters used to calculate the masses of DHTs are derived from the 13 independent parameters already listed in Table 1. which are all obtained by fitting experimental data of the observed hadron mass spectra. This well ensures the correlation between the theoretical predictions of DHT masses and the experimental data of baryon mass spectra, thereby enhancing the reliability of the predictions in this work.

### v2

@@ II.2 Mass formula of DHT in CQM
The masses of the heavy diquarks are determined to be $m_{[cc]^{\bar{3}}_{1}}=2m^{b}_{c}+(-2/3)(B_{cc}+\alpha_{cc}/(m^{b}_{c})^{2})=3300.8$ MeV, $m_{[bb]^{\bar{3}}_{1}}=2m^{b}_{b}+(-2/3)(B_{bb}+\alpha_{bb}/(m^{b}_{b})^{2})=9821.0$ MeV, $m_{[cb]^{\bar{3}}_{0}}=m^{b}_{c}+m^{b}_{b}+(-2/3)(B_{cb}-3\alpha_{cb}/(m^{b}_{b}\cdot m^{b}_{c}))=6567.0$ MeV and $m_{[cb]^{\bar{3}}_{1}}=m^{b}_{c}+m^{b}_{b}+(-2/3)(B_{cb}+\alpha_{cb}/(m^{b}_{b}\cdot m^{b}_{c}))=6600.9$ MeV, respectively. After determining the mass of the heavy diquark $m_{[QQ^{\prime}]^{C}_{S}}$, we need the binding and spin hyperfine coupling constant between $QQ^{\prime}]^{C}_{S}$} and the light quark to calculate the masses of DHTs with Eq. (2). In Refs. Karliner et al. (2009); Karliner and Rosner (2014); Wu and Ma (2023), the studies indicate that there are no bindings between light quark pair ($qq^{\prime}$) and heavy-unflavored quark pair $Qn$. Since the color and spin operators have been explicitly introduced in Eq. (2), the binding between heavy diquark and light quark should be independent of both color and spin, denoted as $B_{[QQ^{\prime}]q}$. So we have $B_{[QQ^{\prime}]q}=B_{Qq}+B_{Q^{\prime}q}$, resulting in $B_{[QQ^{\prime}]n}=0$, $B_{[cc]s}=2B_{cs}=106.8$ MeV, $B_{[bb]s}=2B_{bs}=125.2$ MeV, and $B_{[cb]s}=B_{cs}+B_{bs}=116.0$ MeV. For the spin hyperfine coupling constant $\alpha_{ij}$, the constant is the same for the light-light quarks and heavy-light quarks but is flavor dependent for the heavy-heavy quarks. For $\alpha_{[QQ^{\prime}]^{C}_{S}q}$, since the heavy diquark is regarded as a heavy quark $X$ with HADS, $\alpha_{[QQ^{\prime}]^{C}_{S}q}$ should be $\alpha_{Xq}$, which is proved to be independent on the mass of heavy quark $X$ and share the same value as $\alpha_{qq^{\prime}}$ (abbreviated as $\alpha$) . For detailed discussions, refer to Refs. Karliner et al. (2009); Karliner and Rosner (2014); Wu and Ma (2023).
It should be noted that the hyperfine splitting constant $\alpha_{ij}$ is model-dependent Luo et al. (2017); Zhang et al. (2025); Wu and Ma (2023), while the hyperfine splitting $\alpha_{ij}/{m_{i}m_{j}}$, combined with the spin factor, reflects the mass splitting of different spin configurations of DHTs. According to HADS, the spin splitting should approach zero in the heavy quark limit. This is reflected in Table I, where the spin splitting decreases as the heavy quark mass increases ($|\alpha_{cc}/m_{c}^{2}|>|\alpha_{cb}/m_{c}m_{b}|>|\alpha_{bb}/m_{b}^{2}|$). The impact of HADS symmetry breaking is also worth discussing. In this work, HADS only affects the hyperfine splitting between the heavy diquark and light quark. We therefore define $\chi_{\mathrm{HADS}}=\left|\mathcal{A}_{\mathrm{HADS}}(F_{[QQ^{\prime}]^{C}_{S}}\cdot F_{q})({\sigma}_{[QQ^{\prime}]^{C}_{S}}\cdot{\sigma}_{q}){\alpha_{[QQ^{\prime}]^{C}_{S}q}}/{m_{q}m_{[QQ^{\prime}]^{C}_{S}}}\right|$, where $\mathcal{A}_{\mathrm{HADS}}$ is defined as the degree of HADS breaking. Considering 25% breaking of HADS Hu and Mehen (2006); Wu and Ma (2023), the uncertainties of HADS $\chi_{\mathrm{HADS}}$ in this work are estimated to be at the level of 11.3 MeV, 5.7 MeV and 3.8 MeV for the doubly charmed, bottom-charmed and doubly bottomed tetraquarks, respectively.
The masses, bindings and spin hyperfine splittings needed for calculating the masses of DHTs are listed in Table 2. It should be noted that no new parameters are introduced here. All parameters used to calculate the masses of DHTs are derived from the 13 independent parameters already listed in Table 1. which are all obtained by fitting experimental data of the observed hadron mass spectra. This well ensures the correlation between the theoretical predictions of DHT masses and the experimental data of baryon mass spectra, thereby enhancing the reliability of the predictions in this work.
Table 3: Doubly heavy tetraquark ground states aroused from heavy anti-baryons $\bar{\Lambda}_{c}$($\bar{\Lambda}_{b}$) and $\bar{\Sigma}_{c}$($\bar{\Sigma}_{b}$) with HADS.
State Configuration $I(J^{P})$ Mass formula
$T_{cc}(3876)$ $\bar{n}\bar{n}]^{3}_{0}[cc]^{\bar{3}}_{1}]^{0}_{1}$ $0(1^{+})$ $2m_{n}^{b}+m_{[cc]^{\bar{3}}_{1}}+2\alpha/(m_{n}^{b})^{2}$
$T_{cc}(4035)$ $\bar{n}\bar{n}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{1}_{0}$ $1(0^{+})$ $2m_{n}^{b}+m_{[cc]^{\bar{3}}_{1}}-\frac{2}{3}[\alpha/(m_{n}^{b})^{2}-8\alpha_{[cc]^{\bar{3}}_{1}n}/(m_{n}^{b}m_{[cc]^{\bar{3}}_{1}})]$
$T_{cc}(4058)$ $\bar{n}\bar{n}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{1}_{1}$ $1(1^{+})$ $2m_{n}^{b}+m_{[cc]^{\bar{3}}_{1}}-\frac{2}{3}[\alpha/(m_{n}^{b})^{2}-4\alpha_{[cc]^{\bar{3}}_{1}n}/(m_{n}^{b}m_{[cc]^{\bar{3}}_{1}})]$
$T_{cc}(4103)$ $\bar{n}\bar{n}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{1}_{2}$ $1(2^{+})$ $2m_{n}^{b}+m_{[cc]^{\bar{3}}_{1}}-\frac{2}{3}[\alpha/(m_{n}^{b})^{2}+4\alpha_{[cc]^{\bar{3}}_{1}n}/(m_{n}^{b}m_{[cc]^{\bar{3}}_{1}})]$
$T_{bb}(10396)$ $\bar{n}\bar{n}]^{3}_{0}[bb]^{\bar{3}}_{1}]^{0}_{1}$ $0(1^{+})$ $2m_{n}^{b}+m_{[bb]^{\bar{3}}_{1}}+2\alpha/(m_{n}^{b})^{2}$
$T_{bb}(10586)$ $\bar{n}\bar{n}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{1}_{0}$ $1(0^{+})$ $2m_{n}^{b}+m_{[bb]^{\bar{3}}_{1}}-\frac{2}{3}[\alpha/(m_{n}^{b})^{2}-8\alpha_{[bb]^{\bar{3}}_{1}n}/(m_{n}^{b}m_{[bb]^{\bar{3}}_{1}})]$
$T_{bb}(10593)$ $\bar{n}\bar{n}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{1}_{1}$ $1(1^{+})$ $2m_{n}^{b}+m_{[bb]^{\bar{3}}_{1}}-\frac{2}{3}[\alpha/(m_{n}^{b})^{2}-4\alpha_{[bb]^{\bar{3}}_{1}n}/(m_{n}^{b}m_{[bb]^{\bar{3}}_{1}})]$
$T_{bb}(10608)$ $\bar{n}\bar{n}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{1}_{2}$ $1(2^{+})$ $2m_{n}^{b}+m_{[bb]^{\bar{3}}_{1}}-\frac{2}{3}[\alpha/(m_{n}^{b})^{2}+4\alpha_{[bb]^{\bar{3}}_{1}n}/(m_{n}^{b}m_{[bb]^{\bar{3}}_{1}})]$
$T_{cb}(7142)$ $\bar{n}\bar{n}]^{3}_{0}[cb]^{\bar{3}}_{0}]^{0}_{0}$ $0(0^{+})$ $2m^{b}_{n}+m_{[cb]^{\bar{3}}_{0}}+2\alpha/(m^{b}_{n})^{2}$
$T_{cb}(7176)$ $\bar{n}\bar{n}]^{3}_{0}[cb]^{\bar{3}}_{1}]^{0}_{1}$ $0(1^{+})$ $2m^{b}_{n}+m_{[cb]^{\bar{3}}_{1}}+2\alpha/(m^{b}_{n})^{2}$
$T_{cb}(7347)$ $\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{0}]^{1}_{1}$ $1(1^{+})$ $2m^{b}_{n}+m_{[cb]^{\bar{3}}_{0}}-(2/3)[\alpha/(m^{b}_{n})^{2}]$
$T_{cb}(7358)$ $\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{1}_{0}$ $1(0^{+})$ $2m^{b}_{n}+m_{[cb]^{\bar{3}}_{1}}-(2/3)[\alpha/(m^{b}_{n})^{2}-8\alpha/(m^{b}_{n}m_{[cb]^{\bar{3}}_{1}})]$
$T_{cb}(7369)$ $\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{1}_{1}$ $1(1^{+})$ $2m^{b}_{n}+m_{[cb]^{\bar{3}}_{1}}-(2/3)[\alpha/(m^{b}_{n})^{2}-4\alpha/(m^{b}_{n}m_{[cb]^{\bar{3}}_{1}})]$
$T_{cb}(7392)$ $\bar{n}\bar{n}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{1}_{2}$ $1(2^{+})$ $2m^{b}_{n}+m_{[cb]^{\bar{3}}_{1}}-(2/3)[\alpha/(m^{b}_{n})^{2}+4\alpha/(m^{b}_{n}m_{[cb]^{\bar{3}}_{1}})]$
Table 4: Doubly heavy tetraquark ground states aroused from heavy anti-baryons $\bar{\Omega}_{c}$($\bar{\Omega}_{b}$) with HADS.
State Configuration $J^{P}$ Mass formula
$T_{cc\bar{s}\bar{s}}(4224)$ $\bar{s}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]_{0}$ $0^{+}$ $2m^{b}_{s}+m_{[cc]^{\bar{3}}_{1}}-(2/3)[2B_{[cc]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{s})^{2}-8\alpha/(m^{b}_{s}m_{[cc]^{\bar{3}}_{1}})]$
$T_{cc\bar{s}\bar{s}}(4239)$ $\bar{s}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]_{1}$ $1^{+}$ $2m^{b}_{s}+m_{[cc]^{\bar{3}}_{1}}-(2/3)[2B_{[cc]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{s})^{2}-4\alpha/(m^{b}_{s}m_{[cc]^{\bar{3}}_{1}})]$
$T_{cc\bar{s}\bar{s}}(4270)$ $\bar{s}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]_{2}$ $2^{+}$ $2m^{b}_{s}+m_{[cc]^{\bar{3}}_{1}}-(2/3)[2B_{[cc]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{s})^{2}+4\alpha/(m^{b}_{s}m_{[cc]^{\bar{3}}_{1}})]$
$T_{bb\bar{s}\bar{s}}(10740)$ $\bar{s}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]_{0}$ $0^{+}$ $2m^{b}_{s}+m_{[bb]^{\bar{3}}_{1}}-(2/3)[2B_{[bb]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{s})^{2}-8\alpha/(m^{b}_{s}m_{[bb]^{\bar{3}}_{1}})]$
$T_{bb\bar{s}\bar{s}}(10745)$ $\bar{s}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]_{1}$ $1^{+}$ $2m^{b}_{s}+m_{[bb]^{\bar{3}}_{1}}-(2/3)[2B_{[bb]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{s})^{2}-4\alpha/(m^{b}_{s}m_{[bb]^{\bar{3}}_{1}})]$
$T_{bb\bar{s}\bar{s}}(10755)$ $\bar{s}\bar{s}]^{3}_{1}[[bb]^{\bar{3}}_{1}]_{2}$ $2^{+}$ $2m^{b}_{s}+m_{[bb]^{\bar{3}}_{1}}-(2/3)[2B_{[bb]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{s})^{2}+4\alpha/(m^{b}_{s}m_{[bb]^{\bar{3}}_{1}})]$
$T_{cb\bar{s}\bar{s}}(7508)$ $\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{0}]_{1}$ $1^{+}$ $2m^{b}_{s}+m_{[cb]^{\bar{3}}_{0}}-(2/3)[2B_{[cb]^{\bar{3}}_{0}s}+\alpha/(m^{b}_{s})^{2}]$
$T_{cb\bar{s}\bar{s}}(7527)$ $\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]_{0}$ $0^{+}$ $2m^{b}_{s}+m_{[cb]^{\bar{3}}_{1}}-(2/3)[2B_{[cb]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{s})^{2}-8\alpha/(m^{b}_{s}m_{[cb]^{\bar{3}}_{1}})]$
$T_{cb\bar{s}\bar{s}}(7535)$ $\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]_{1}$ $1^{+}$ $2m^{b}_{s}+m_{[cb]^{\bar{3}}_{1}}-(2/3)[2B_{[cb]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{s})^{2}-4\alpha/(m^{b}_{s}m_{[cb]^{\bar{3}}_{1}})]$
$T_{cb\bar{s}\bar{s}}(7550)$ $\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]_{2}$ $2^{+}$ $2m^{b}_{s}+m_{[cb]^{\bar{3}}_{1}}-(2/3)[2B_{[cb]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{s})^{2}+4\alpha/(m^{b}_{s}m_{[cb]^{\bar{3}}_{1}})]$
Table 5: Doubly heavy tetraquark ground states aroused from heavy anti-baryons $\bar{\Xi}_{c}$($\bar{\Xi}_{b}$) with HADS.
State Configuration $I(J^{P})$ Mass formula
$T_{cc\bar{s}}(4026)$ $\bar{n}\bar{s}]^{3}_{0}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[cc]^{\bar{3}}_{1}}-(2/3)[B_{[cc]^{\bar{3}}_{1}s}-3\alpha/(m^{b}_{n}m^{b}_{s})]$
$T_{cc\bar{s}}(4127)$ $\bar{n}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{0}$ $\frac{1}{2}(0^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[cc]^{\bar{3}}_{1}}-(2/3)[B_{[cc]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{n}m^{b}_{s})-4\alpha/(m^{b}_{n}m_{[cc]^{\bar{3}}_{1}})-4\alpha/(m^{b}_{s}m_{[cc]^{\bar{3}}_{1}})]$
$T_{cc\bar{s}}(4146)$ $\bar{n}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[cc]^{\bar{3}}_{1}}-(2/3)[B_{[cc]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{n}m^{b}_{s})-2\alpha/(m^{b}_{n}m_{[cc]^{\bar{3}}_{1}})-2\alpha/(m^{b}_{s}m_{[cc]^{\bar{3}}_{1}})]$
$T_{cc\bar{s}}(4184)$ $\bar{n}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]^{\frac{1}{2}}_{2}$ $\frac{1}{2}(2^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[cc]^{\bar{3}}_{1}}-(2/3)[B_{[cc]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{n}m^{b}_{s})+2\alpha/(m^{b}_{n}m_{[cc]^{\bar{3}}_{1}})+2\alpha/(m^{b}_{s}m_{[cc]^{\bar{3}}_{1}})]$
$T_{bb\bar{s}}(10534)$ $\bar{n}\bar{s}]^{3}_{0}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[bb]^{\bar{3}}_{1}}-(2/3)[B_{[bb]^{\bar{3}}_{1}s}-3\alpha/(m^{b}_{n}m^{b}_{s})]$
$T_{bb\bar{s}}(10660)$ $\bar{n}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{0}$ $\frac{1}{2}(0^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[bb]^{\bar{3}}_{1}}-(2/3)[B_{[bb]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{n}m^{b}_{s})-4\alpha/(m^{b}_{n}m_{[bb]^{\bar{3}}_{1}})-4\alpha/(m^{b}_{s}m_{[bb]^{\bar{3}}_{1}})]$
$T_{bb\bar{s}}(10666)$ $\bar{n}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[bb]^{\bar{3}}_{1}}-(2/3)[B_{[bb]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{n}m^{b}_{s})-2\alpha/(m^{b}_{n}m_{[bb]^{\bar{3}}_{1}})-2\alpha/(m^{b}_{s}m_{[bb]^{\bar{3}}_{1}})]$
$T_{bb\bar{s}}(10679)$ $\bar{n}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{2}$ $\frac{1}{2}(2^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[bb]^{\bar{3}}_{1}}-(2/3)[B_{[bb]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{n}m^{b}_{s})+2\alpha/(m^{b}_{n}m_{[bb]^{\bar{3}}_{1}})+2\alpha/(m^{b}_{s}m_{[bb]^{\bar{3}}_{1}})]$
$T_{cb\bar{s}}(7286)$ $\bar{n}\bar{s}]^{3}_{0}[cb]^{\bar{3}}_{0}]^{\frac{1}{2}}_{0}$ $\frac{1}{2}(0^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[cb]^{\bar{3}}_{0}}-(2/3)[B_{[cb]^{\bar{3}}_{0}s}-3\alpha/(m^{b}_{n}m^{b}_{s})]$
$T_{cb\bar{s}}(7320)$ $\bar{n}\bar{s}]^{3}_{0}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[cb]^{\bar{3}}_{1}}-(2/3)[B_{[cb]^{\bar{3}}_{1}s}-3\alpha/(m^{b}_{n}m^{b}_{s})]$
$T_{cb\bar{s}}(7425)$ $\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{0}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[cb]^{\bar{3}}_{0}}-(2/3)[B_{[cb]^{\bar{3}}_{0}s}+\alpha/(m^{b}_{n}m^{b}_{s})]$
$T_{cb\bar{s}}(7440)$ $\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{0}$ $\frac{1}{2}(0^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[cb]^{\bar{3}}_{1}}-(2/3)[B_{[cb]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{n}m^{b}_{s})-4\alpha/(m^{b}_{n}m_{[cb]^{\bar{3}}_{1}})-4\alpha/(m^{b}_{s}m_{[cb]^{\bar{3}}_{1}})]$
$T_{cb\bar{s}}(7449)$ $\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{1}$ $\frac{1}{2}(1^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[cb]^{\bar{3}}_{1}}-(2/3)[B_{[cb]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{n}m^{b}_{s})-2\alpha/(m^{b}_{n}m^{b}_{[cb]^{\bar{3}}_{1}})-2\alpha/(m^{b}_{s}m_{[cb]^{\bar{3}}_{1}})]$
$T_{cb\bar{s}}(7468)$ $\bar{n}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]^{\frac{1}{2}}_{2}$ $\frac{1}{2}(2^{+})$ $m^{b}_{n}+m^{b}_{s}+m_{[cb]^{\bar{3}}_{1}}-(2/3)[B_{[cb]^{\bar{3}}_{1}s}+\alpha/(m^{b}_{n}m^{b}_{s})+2\alpha/(m^{b}_{n}m_{[cb]^{\bar{3}}_{1}})+2\alpha/(m^{b}_{s}m_{[cb]^{\bar{3}}_{1}})]$